# Safe Control using Occupancy Grid Map-based Control Barrier Function (OGM-CBF)

### v2

@@ Abstract
Safe navigation in unknown environments stands as a significant challenge in the field of robotics. Control Barrier Function (CBF) is a strong mathematical tool to guarantee safety requirements. However, a common assumption in many works is that the CBF is already known and obstacles have predefined shapes. In this letter, we present a novel method called Occupancy Grid Map-based Control Barrier Function (OGM-CBF), which defines Control Barrier Function based on Occupancy Grid Maps. This enables generalization to unknown environments while generating online local or global maps of the environment using onboard perception sensors such as LiDAR or camera. With this method, the system guarantees safety via a single, continuously differentiable CBF per time step, which can be represented as one constraint in the CBF-QP optimization formulation while having an arbitrary number of obstacles with unknown shapes in the environment. This enables practical real-time implementation of CBF in both unknown and known environments. The efficacy of OGM-CBF is demonstrated in the safe control of an autonomous car in the CARLA simulator and a real-world industrial mobile robot.
@@
Autonomous Vehicle Navigation, Collision Avoidance, Vision-Based Navigation, Sensor-based Control.
@@ I INTRODUCTION
Applications of mobile robots in domains such as autonomous vehicles, industrial robots, and multi-agent systems are often tasked to respect safety while navigating in complex and unknown environments, which stands as one of the most formidable challenges in the field of robotics [1]. One promising approach to integrate safety into design is via Control Barrier Function (CBF)[2] due to the strong mathematical guarantee of satisfying safety requirements. CBFs have been applied in various fields such as legged robots [3], autonomous vehicles [4], multi-robot systems [5] [6], decentralized multi-agent control [7], and drones [8].
In recent years, CBF has been successfully integrated with Control Lyapunov Function (CLF) via Quadratic Programs (QP) for safety critical systems [9], enabling the simultaneous satisfaction of safety and stability. Moreover, given a potentially unsafe nominal control policy, CBF can be embedded as a QP-based safety filter to modify the unsafe control command in a minimally invasive manner to prevent violations of the safety constraints[2].
@@
In designing CBF, it is generally assumed that unsafe sets are already known (usually as circular shapes), which is not applicable to unsafe sets with arbitrary shapes in uncharted environments. Practical implementation of CBF in unknown environments relies on a sufficient awareness of the environment, unsafe regions, and their changes to reach a quick, yet smooth generation of a safety certificate as a response. This requires online estimation of unsafe sets and their integration into the CBF formulation.
One practical approach to operate safely in unknown and dynamic environments without sufficient prior information about the environment is the integration of sensory data into the CBF formulation. This enables the online synthesis of CBF based on the aggregated sensor data from perceptive sensors such as LiDAR or camera.
To construct CBF based on sensor data, one approach is to learn both the classification of safe and unsafe regions and CBF simultaneously based on sensor measurements. In [10], the authors used neural networks to learn the CBF directly in the observation space using local observations from LiDAR, without the need for the robot’s state estimation and localization. In [11], the authors designed a neural network, called BarrierNet, to learn CBF with respect to the user-defined safe sets, and in [12], a neural network was trained to extract features from front-view images and learn the BarrierNet as an end-to-end vision-based solution for lane keeping and obstacle avoidance. In [13], CBF is synthesized based on the online approximation of the signed distance function from LiDAR measurements using an incremental training approach with replay memory, and in [14], CBF is constructed based on the classification of safe and unsafe sets using a support vector machine on LiDAR measurements. In these approaches, sensor data at each time step is immediately employed to construct CBF, resulting in a lack of historical memory for obstacles that have been previously seen. Moreover, these methods are bound to specific sensor modalities and cannot combine information from different sensors for CBF generation.
One other approach is to distinguish the tasks of classifying safe and unsafe areas from the synthesis of CBF, treating them as separate modules. Consequently, established classification techniques, such as semantic segmentation and mapping, can be utilized to identify safe and unsafe areas, upon which CBF can be defined. This modular approach can enhance the efficacy of the final controller, particularly in the case of significant changes in the type of environment where the safe controller is deployed. In V-CBF [15], Control Barrier Functions are defined based on segmented unsafe sets using RGBD cameras. Although this framework successfully navigates safely and avoids obstacles with unknown shapes in uncharted environments, it relies solely on instantaneous sensor inputs and lacks memory of previously encountered obstacles. Moreover, as the robot is considered as a point, to cover the size and shape of the robot, a handful of the most conservative points of the constructed CBF should be selected and added as multiple constraints in QP formulation for a successful safe navigation. It is not trivial how these points are selected.
One practical way to approximate the shapes of obstacles is Occupancy Grid Map (OGM). OGM, as a multi-dimensional discrete map, efficiently captures the complexity of the surrounding environment, including the state and shape of all detected obstacles. With accurate state estimation for the robot, the OGM illustrates the relative positions of obstacles with respect to the robot, providing comprehensive situational awareness. The map employs a probabilistic formulation to store all previously seen obstacles, retaining information about these past encounters. Additionally, obstacles within the OGM can be inflated according to the robot’s size, enhancing safe navigation. OGM not only embeds essential information for the robot’s safe navigation, including the position and inflated shape of obstacles, but it also can be constructed online using real-time sensor data. Online mapping facilitates the robot’s navigation in unknown environments since mapping and synthesizing CBFs can be done in real time. Moreover, various sensor configurations can be utilized to generate OGMs, making it a versatile tool for representing unsafe zones. However, OGM alone cannot be directly used for generating CBF since OGM behaves in a binary manner and does not hold correct variational behavior near the boundary of unsafe sets. Thus, a domain translation function is necessary to transform the occupancy grid map into a representation suitable for CBF synthesis. In [16], CBF was designed using elliptical outer approximation on occupancy grid maps. However, the obstacles were approximated with minimum bounding circles, which can be overly conservative. Additionally, this approach requires a separate constraint for each obstacle.
In this paper, we introduce a novel method to define CBF based on occupancy grid maps called Occupancy Grid Map-based Control Barrier Function (OGM-CBF). Our contribution is as follows:
  • CBF is represented as a single constraint at each time step in a QP formulation while assuming the presence of an arbitrary number of obstacles with unknown shapes, enabling real-time and computationally light optimization to guarantee safety.
  • The designed safe controller is sensor agnostic. In addition, CBF can be constructed from hybrid sensor modalities using occupancy grid maps.
  • The constructed CBF has a memory of all previously seen obstacles.
  • The constructed CBF accounts for the size and shape of the robot using an inflated occupancy grid map.
@@
  • We evaluated the efficacy of the proposed method in the CARLA simulator in complex scenarios as well as with an industrial mobile robot in real-world experiments.
@@
The outline of this paper is as follows. Section II reviews the principles of Control Barrier Function, Occupancy Grid Map, and Signed Distance Function. Section III describes our designed control methodology. Section IV explains the designed smoothing function used in our CBF. Section V is dedicated to the demonstration of the introduced control strategy in the CARLA simulator and on a real robot. Finally, in section VI, the conclusions and future work are provided.
@@ II PRELIMINARIES AND BACKGROUND
$z$ $=O(\boldsymbol{x},\boldsymbol{q})$
where $\boldsymbol{x}\in\mathbb{X}\subseteq\mathbb{R}^{n}$, $f:\mathbb{R}^{n}\to\mathbb{R}^{n}\quad\text{and}\quad g:\mathbb{R}^{n}\to\mathbb{R}^{2}$ are locally Lipschitz continuous functions, $O:\mathbb{R}^{n}\times\mathbb{R}^{q}\rightarrow\mathbb{R}^{k}$ is a perception sensor module such as camera or LiDAR capable of generating occupancy grid maps, $\boldsymbol{x}$ and $z$ are the state and observation vectors, respectively, and $\boldsymbol{q}$ is environment variables. In this paper, our system is a kinematic model of a robot with $\boldsymbol{u}=[\mathrm{v},\boldsymbol{\omega}]^{T}$, the linear and angular speeds as input.
@@
Let $f:D\subseteq\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a function. $f$ is locally Lipschitz continuous at a point $x\in D$ if there exist constants $M>0$ and $\delta>0$ such that for all $x^{\prime}\in D$ with $\lVert x-x^{\prime}\rVert\leq\delta$, the inequality $\lVert f(x)-f(x^{\prime})\rVert\leq M\lVert x-x^{\prime}\rVert$ is satisfied.
@@ II-A Control Barrier Function
$S=\left\{\boldsymbol{x}\in\mathbb{X}\;|\;h(\boldsymbol{x})\geq 0\right\}$ (2)
To keep the system (1) remain in the safe set for all time, the set $S$ should be a forward invariant set.
@@
As proven in [2], given the dynamic system (1), the function $h(\boldsymbol{x})$ is a control barrier function if there exists an extended class $\kappa$ function $\alpha(.)$ such that for all $\boldsymbol{x}\in\mathbb{X}$:
@@ II-B Occupancy Grid Map
Occupancy grid maps (OGMs) are fundamental tools in mobile robotics for mapping and navigation. They represent the environment as a tessellated, multi-dimensional grid, with each cell quantifying whether a part of the environment is occupied or not [18]. This setup, given the state of the system, incorporates a probabilistic approach to account for uncertainties in measurements and the robot’s pose estimation [19], [20]. OGMs are typically updated using sensor data from LiDAR, sonar, or vision systems and the robot’s movements over time. This process can be employed using various frameworks, including probabilistic [18], Bayesian [21], neural network-based [22], and forward modeling approaches [23].
The process of dynamically generating an occupancy grid map based on the current robot state and sensor inputs can be succinctly described as follows:
$m^{\prime}=M(\boldsymbol{x},z,m),$ (4)
where $\boldsymbol{x}$ represents the robot’s current state, $z$ denotes the measurements from onboard perception sensors, and $m$ is the existing map. The function $M:\mathbb{R}^{n}\times\mathbb{R}^{k}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ updates the current map belief $m$ to $m^{\prime}$, by integrating new sensor data at the latest state of the robot into the map.
This formulation reflects that the occupancy grid map continuously evolves, capturing real-time changes in the environment as detected by the sensors and the movement of the robot.
In this paper, the occupancy grid map $m$ is converted into a binary representation. In the binary OGM, each occupied cell signifies an obstacle in the environment. Consequently, the overall unsafe set, denoted by $\mathbb{X}_{u}$, is defined as the union of all states corresponding to the occupied cells of the map. Trivially, the final safe set $\mathbb{X}_{s}$ is given by $\mathbb{X}\setminus\mathbb{X}_{u}$, which includes all states where the robot is located in non-occupied cells of the map. The process of deriving the safe set $\mathbb{X}_{s}$ from map $m$ can be described briefly as follows:
$\mathbb{X}_{s}=M_{b}(m)$ (5)
@@ II-C Signed Distance Function
A signed distance function (SDF) is a function that, for a given point $p$, outputs the signed distance from $p$ to the closest boundary of a set $\Omega$ in a metric space. The sign of the distance indicates whether $p$ is inside or outside $\Omega$ [24]. In this paper, this function is defined to take positive values for points inside $\Omega$, decrease in value as $p$ approaches the boundary where the distance is zero, and take negative values outside $\Omega$. If $\Omega$ is a subset of a metric space with metric $d$, the described signed distance function is defined as follows:
$\phi(p,\Omega)\equiv$ $d(p,\partial\Omega)$ if $p\in\Omega$ (6a)
$\phi(p,\Omega)\equiv$ $-d(p,\partial\Omega)$ if $p\notin\Omega$ (6b)
where $\partial\Omega$ denotes the boundary of the set $\Omega$ and $d(p,\partial\Omega)$ represents the minimum distance from $p$ to points on the boundary. Additionally, $d(.,\partial\Omega)$ is the desired distance metric to the set $\partial\Omega$.
In the case where $\Omega$ is a subset of Euclidean space $\mathbb{R}^{n}$ with a piecewise smooth boundary, and $d(.,\partial\Omega)$ is the Euclidean distance metric, the signed distance function is differentiable almost everywhere. Moreover, the gradient of the function satisfies the eikonal equation $|\nabla\phi|=1$ at every point where it is differentiable [25].
@@ III CONTROL METHODOLOGY
In this section, we introduce a novel method for defining CBF using occupancy grid maps. Like traditional CBFs, the designed CBF can be applied using various formulations, such as a safety filter or a CBF-CLF QP. For our experimental evaluations, we have chosen the CBF-CLF QP formulation. Essential details of this formulation are succinctly provided in this section.

### v3

# Safe Robot Control using Occupancy Grid Map-based Control Barrier Function (OGM-CBF)

@@ Abstract
Safe control in unknown environments is a significant challenge in robotics. While Control Barrier Functions (CBFs) are widely used to guarantee system safety, they often assume known environments with predefined obstacles. The proposed method constructs CBFs directly from perception sensor input and introduces a new first-order barrier function for a 3D kinematic robot motion model. The proposed CBF is constructed by combining Occupancy Grid Mapping (OGM) and Signed Distance Functions (SDF). The OGM framework abstracts sensor inputs, making the solution compatible with any sensor modality capable of generating occupancy maps. Moreover, the OGM enhances situational awareness along the robot’s motion trajectory, by integrating both current and previously mapped data. The SDF encapsulates complex obstacle shapes defined by OGM into real-time computable values, enabling the method to handle obstacles of arbitrary shapes. This enables a single constraint in the CBF-QP optimization for each point on the robot, regardless of the number or shape of obstacles. The effectiveness of the proposed approach is demonstrated through simulations on autonomous driving in the CARLA simulator and real-world experiments with an industrial mobile robot, using a simplified 2D version of the method.
@@
Sensor-based Control, Collision Avoidance, Vision-Based Navigation, Autonomous Vehicle Navigation.
@@ I INTRODUCTION
Applications of mobile robots in domains such as autonomous vehicles, industrial robots, and multi-agent systems are often tasked to respect safety while navigating in complex and unknown environments, which stands as one of the most formidable challenges in the field of robotics [1]. One promising approach to integrate safety into design is via Control Barrier Function (CBF)[2] due to the strong mathematical guarantee of satisfying safety requirements. CBFs have been applied in various fields such as legged robots [3], autonomous vehicles [4], multi-robot systems [5], and drones [6]. In recent years, CBF has been combined with Control Lyapunov Functions (CLF) in Quadratic Programs (QP) to ensure both safety and stability in safety-critical systems [7]. Additionally, CBF can act as a QP-based safety filter, modifying unsafe control commands minimally to prevent safety violations [2].
@@
In designing CBF, it is generally assumed that unsafe sets are already known (usually as circular shapes), which is not applicable to unsafe sets with arbitrary shapes in uncharted environments. Practical implementation of CBF in unknown environments relies on a sufficient awareness of the environment, unsafe regions, and their changes to reach a quick, yet smooth generation of a safety certificate as a response. This requires online estimation of unsafe sets using onboard sensory data and their integration into the CBF formulation as well as advanced frameworks for tuning CBF parameters. For example, in [8], the authors design a neural network, called BarrierNet, to optimize parameters of CBF constructed from sensory data with respect to the user-defined safe sets. Constructing CBF from onboard sensors has been addressed in the literature in through multiple approaches.
One common approach involves detecting the state of obstacles using onboard measurements and then fitting classical circular shapes to the obstacles. This defines the CBF based on the robot’s state, the obstacle’s state, and an approximated radius using traditional methods. In [8], a neural network model was trained to extract the state of obstacles and the robot from front-view images and construct high order CBFs, while tuning their parameters through BarrierNet. The constructed CBFs in this method are dependent on position of obstacles, while simplifying the shape of obstacles via disk fitting. Consequently, using this circular formulation for CBFs, each obstacle requires a separate BarrierNet neuron and constraint, with the number of obstacles needing to be known in advance. In [9], CBF was designed using elliptical outer approximation on occupancy grid maps. The approximation of obstacles with minimum bounding circles in these methods can be overly conservative, and, similarly, this approach requires a separate constraint for each obstacle.
Another approach involves using sensor measurements, such as LiDAR, to approximate signed distance functions (SDFs) and construct CBFs based on the approximated SDF. SDFs provide a smooth, continuous representation of the closest distance to a set’s surface, offering essential information for online obstacle avoidance and motion planning [10]. In [11], safe and unsafe sets are classified using a support vector machine on LiDAR measurements, and the CBF is constructed as the signed distance function of the recovered unsafe set. In this method, the CBF is expressed as a single constraint in the QP optimization, capturing the entire unsafe set. However, the online construction of the unsafe set while accounting for previously observed obstacles requires an exhaustive data collection process, which can become computationally expensive. To address this issue, [12] approximated SDF from LiDAR measurements using an incremental training approach with replay memory. In this approach, the online training set includes the latest observations and a randomly sampled subset of previous observations. However, SDF is approximated separately for each obstacle, and consequently, one CBF constraint is needed per obstacle. It is important to note that both of these methods rely on a single sensor modality, LiDAR sensors.
CBF synthesis can also be achieved by classifying safe and unsafe sets in the observation space and defining CBFs directly within that space. In V-CBF [13], semantic segmentation in the image frame is used to capture the shape of obstacles using RGB-D cameras, and a domain translation is applied using cGAN networks to transform the segmented unsafe set into a smooth domain where the CBF is defined. Although this framework successfully enables safe navigation and obstacle avoidance of unknown shapes in uncharted environments, it relies solely on instantaneous sensor inputs from RGB-D cameras and lacks memory of previously encountered obstacles.
In this paper, we introduce a novel method called OGM-CBF, which defines CBF using OGM and SDF. Our contributions are as follows:
  • A novel method to construct CBF from perception sensors using Occupancy Grid Mapping (OGM) and Signed Distance Functions (SDF) for robot 3D kinematic motion, and the corresponding novel barrier function leading to first order CBF.
  • The OGM component abstracts the sensor input, enabling our method to work with any sensor modality that can generate OGM, a map that safe and unsafe cells are classified.
  • The OGM component adds memory enabling our method to integrate sensor data along the motion trajectory increasing situational awareness compared to instantaneous sensor input. Naturally, previously mapped information can also be used.
  • In our formulation, SDF encapsulates complex unsafe sets defined by OGM to few numbers that can be computed in real-time, enabling our method to work with obstacles of arbitrary shapes. Robot shape is resolved by calculating SDF at number of representative points on the robot or inflating OGM, and one constraint per point in a QP problem.
@@
  • We evaluated the efficacy of the proposed method on safe control of an autonomous driving car in the CARLA simulator as well as an industrial mobile robot in real-world experiments, using a simplified 2D version of our method.
@@
The outline of this paper is as follows : Section II reviews the principles of Control Barrier Function, Occupancy Grid Map, and Signed Distance Function. Section III describes our designed control methodology. Section IV provides a qualitative comparison. Section V is dedicated to the demonstration of the introduced control strategy in the CARLA simulator and on a real robot. Finally, in section VI, the conclusions and future work are provided.
@@ II PRELIMINARIES AND BACKGROUND
$z=O_{1}(\boldsymbol{x},\boldsymbol{q})$
$(\boldsymbol{p},\hat{\boldsymbol{x}})=O_{2}(\boldsymbol{x})$
where $\boldsymbol{x}\in\mathbb{X}\subseteq\mathbb{R}^{n}$ is the state of the system, $\boldsymbol{u}=[\boldsymbol{v},\boldsymbol{\omega}]^{T}\in\mathbb{U}\subseteq\mathbb{R}^{m}$ is the control input as linear and angular velocities in body frame, and $f:\mathbb{R}^{n}\to\mathbb{R}^{n}\quad\text{and}\quad g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ are locally Lipschitz continuous functions. We assume our robot is equipped with two observation systems $O_{1}$ and $O_{2}$. Observation $O_{1}:\mathbb{R}^{n}\times\mathbb{R}^{q}\rightarrow\mathbb{R}^{k}$ is a perception sensor capable of generating occupancy grid maps, such as camera, LiDAR, or radar. $O_{1}$ is a function of the state and the environment variables $\boldsymbol{q}\in\mathbb{R}^{q}$, providing the observation vector $z\in\mathbb{R}^{k}$. The function $O_{2}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{i}\times\mathbb{R}^{i}$ is a function of states and provides position $\boldsymbol{p}\in\mathbb{R}^{i}$ and the robot heading unit vector $\hat{\boldsymbol{x}}\in\mathbb{R}^{i}$.
@@
Function $f:D\subseteq\mathbb{R}^{n}\to\mathbb{R}^{m}$ is locally Lipschitz continuous at a point $x\in D$, if there exist constants $M>0$ and $\delta>0$ such that for all $x^{\prime}\in D$ with $\lVert x-x^{\prime}\rVert\leq\delta$, the inequality $\lVert f(x)-f(x^{\prime})\rVert\leq M\lVert x-x^{\prime}\rVert$ is satisfied, with $\lVert$.$\rVert$ . ∥ denoting the L2 norm.
@@ II-A Control Barrier Function
$S=\left\{\boldsymbol{x}\in\mathbb{X}\;|\;h(\boldsymbol{x})\geq 0\right\}.$ (2)
To study the safety of system (1), we will use the concept of forward invariant set defined next.
@@
As defined in [2], given the dynamic system (1), the function $h(\boldsymbol{x})$ is a control barrier function if there exists an extended class $\kappa$ function $\alpha(.)$ such that for all $\boldsymbol{x}\in\mathbb{X}$:
@@ II-B Occupancy Grid Map
Occupancy grid maps (OGMs) are fundamental tools in mobile robotics for mapping and navigation. They represent the environment as a tessellated, multi-dimensional grid, with each cell quantifying whether a part of the environment is occupied or not [15]. OGMs are updated and refined over time to improve environmental representation, often using a probabilistic approach to account for uncertainties in sensor measurements [16]. Moreover, OGM is a sensor-agnostic representation which can be updated using measurements from various sensor modalities such as LiDAR, RGB camera [17], stereo camera [18], and radar [19]. Robot pose trajectory is then used to integrate observations over time, succinctly described as follows.
$m^{\prime}=M(\boldsymbol{p},\hat{\boldsymbol{x}},z,m),$ (4)
where $\boldsymbol{p}$ represents the robot’s position, $\hat{\boldsymbol{x}}$ is the robot’s heading, $z$ denotes the sensor measurements from onboard perception sensors, and $m$ is the current belief about the map. The function $M:\mathbb{R}^{i}\times\mathbb{R}^{i}\times\mathbb{R}^{k}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ updates the map belief $m$ to $m^{\prime}$, by integrating new sensor data at the latest pose of the robot into the map. Notice that we need the robot pose (position and heading) to add sensor observations $z$ to current map $m$ to get $m^{\prime}$ and after generation of the map, map is independent of robots pose and sensor measurement. Sensor extrinsic calibration values (sensor pose in body frame) are used to transform the sensor data into body frame.
To define the safe set, we convert the occupancy grid map $m$ into a binary representation. Equation (5b) explains the process of binarizing the occupancy grid map, where $\mu$ is defined as the minimum probability (or confidence level) above which a grid cell is considered to be occupied by an obstacle thus unsafe. In other words, $\mu$ is the value that determines the acceptance of risk in classifying a grid cell as an obstacle.
$M_{b}(m(\boldsymbol{p}))=$ 00 0 if $m(\boldsymbol{p})<\mu$ (5a)
$M_{b}(m(\boldsymbol{p}))=$ 11 11 if $m(\boldsymbol{p})\geq\mu$ (5b)
Notice that with slight abuse of notation, we use $\boldsymbol{p}$ to indicate a cell in the grid map. We can now define the unsafe set $\mathbb{X}_{u}$ as the union of all states corresponding to the occupied cells of the map, formally,
$\mathbb{X}_{u}=\{\boldsymbol{x}\in\mathbb{X}\mid M_{b}(m(O_{2}(\boldsymbol{x})))=0\}.$ (6)
Consequently, the safe set $\mathbb{X}_{s}$ is given by $\mathbb{X}\setminus\mathbb{X}_{u}$, which includes all the positions in non-occupied cells of the map.
@@ II-C Signed Distance Function
A signed distance function (SDF) is a function that, for a given point $\boldsymbol{p}$, outputs the signed distance from $\boldsymbol{p}$ to the closest point of a set $\Omega$ in a metric space. The sign of the distance indicates whether $\boldsymbol{p}$ is inside or outside $\Omega$ [20]. If $\Omega$ is a subset of a metric space with metric $d$, the signed distance function is defined as follows:
[…]
where $\partial\Omega$ denotes the boundary of the set $\Omega$ and $d(\boldsymbol{p},\partial\Omega)$ represents the minimum distance from $\boldsymbol{p}$ to points on the boundary. Additionally, $d(.,\partial\Omega)$ is the desired distance metric to the set $\partial\Omega$.
In the case where $\Omega$ is a subset of Euclidean space $\mathbb{R}^{n}$ with a piecewise smooth boundary, and $d(.,\partial\Omega)$ is the Euclidean distance metric, the signed distance function is differentiable almost everywhere. Moreover, the gradient of the function satisfies the eikonal equation $|\nabla\phi|=1$ at every point where it is differentiable. Interested readers are invited to see more details in [21], [22].
@@ III CONTROL METHODOLOGY
In this section, we introduce our novel method. Typically, CBF is added as a safety filter to a performance/stabilizing controller. Performance controllers that can be formulated as QP constraints, such as, Control Lyapunov Functions (CLF), can be combined in a QP optimization together with CBF constraints [2]. Formulation of safety controllers and CBF as a QP optimization was an important achievement introduced in [2], paving the way to many proceeding work in safe control, including ours. For our experimental evaluations, we have chosen the CBF-CLF QP formulation. Essential details of this formulation are succinctly provided in this section.